OpenStudy (anonymous):

solve for y: 5x-y= -3

OpenStudy (anonymous):

5x - y = -3 get 5x to the other side by - 5x to both sides <-- can you do this?

OpenStudy (anonymous):

i think i got it. is it y = 3 - 5x

OpenStudy (anonymous):

almost correct lol, but make that - a + 5x - y = -3 -y = -3 - 5x y = 3 + 5x